# Friend or Foe? Harnessing Controllable Overfitting for Anomaly Detection

### v1

@@ 6 Dual Control Mechanism Design Concept
7 Main Functional Pseudocode
@@
1:   Initialize RADI value: $\textit{RADI}\leftarrow 0$
@@
8 Extended Discussion of Gaussian Noise Validation
8.1 Theoretical Rationale for Gaussian Noise
In this work, we chose to use Gaussian noise as a preliminarily pseudo-anomaly generator due to its statistical properties and simplicity. Gaussian noise, characterized by its mean and variance, serves as an ideal candidate for generating random deviations that can mimic unexpected patterns in normal data.
@@ 8.1 Theoretical Rationale for Gaussian Noise
8.2 Validation Through Total Variation Distance (TVD)
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we conducted an empirical analysis using the Total Variation Distance (TVD) metric. TVD is a standard measure used in probability theory to quantify the dissimilarity between two distributions. We calculated TVD to evaluate how closely Gaussian noise could mimic true anomalies in both spatial and intensity domains.
@@ 8.2 Validation Through Total Variation Distance (TVD)
We conducted experiments on our dataset using both generated Gaussian noise and real anomalies. presents a comparison of the TVD values for different classes. The results indicate that the Gaussian distribution closely matches the distribution of real normal and anomalies across a range of classes. Specifically, it can be seen in  Sec. 4.4.
8.3 Experimental Results and Qualitative Study
@@
To further validate the results, we present detailed histograms for the anomaly detection scores obtained across 15 individual categories as well as the overall dataset. These histograms were generated under the condition where the Anomaly Rate Quotient (ARQ) is set to 0.006, using RD++ as our baseline framework. Each histogram in  Figs. 3 and 4 provides a clear visualization of the distribution of detection scores, overlaid with a Gaussian fitting curve for enhanced interpretability. This fitting curve highlights the distribution, allowing for a more effective comparison of the model’s performance across different categories and the overall dataset, further solidifying the reliability of our evaluation metrics.
@@
Figure 3: Histograms for anomaly scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
@@
Figure 4: Histograms for normal scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
8.4 Future Directions with GLASS-like Pseudo-Anomaly Generators
While Gaussian noise has proven effective as a preliminary pseudo-anomaly generator, future work could explore more advanced methods inspired by the GLASS approach [4]. The GLASS method introduces a structured way to generate pseudo-anomalies by leveraging geometric transformations and localized pattern synthesis, providing greater control over anomaly characteristics.
9 Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
To further validate the effectiveness of our methods, we provide a qualitative comparison of anomaly detection results across 15 distinct categories using baselines such as RD, RD++, UniAD, DiAD, and the proposed COAD framework in  Figs. 5, 6, 7 and 8. Each subfigure demonstrates: (from left to right) the original image, the ground truth anomaly mask, the result from the baseline method, and the result after applying COAD.
The comparison shows that COAD effectively reduces both false positive rates (FPR) and false negative rates (FNR), accurately capturing anomaly regions while minimizing misclassifications. These results are consistent with the improvements reflected in the earlier quantitative data in  Tabs. 3 and 4, where COAD consistently outperforms its counterparts in image-level and pixel-level AUROC metrics, highlighting COAD’s capability to outperform the baselines.
@@ 9 Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
Figure 5: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using the Baseline Method RD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 6: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method RD++ and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 7: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method UniAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 8: Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method DiAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
10 Future Work
In future work, we plan to refine our pseudo-anomaly generators by incorporating advanced methods in  Sec. 8.4, enabling the generation of more structured and domain-relevant pseudo-anomalies.

### v2

@@ Appendix A Dual Control Mechanism Design Concept
Appendix B Main Functional Pseudocode
@@
[…]
@@
Appendix C Extended Discussion of Gaussian Noise Validation
C.1. Theoretical Rationale for Gaussian Noise
In this work, we choose to use Gaussian noise as a preliminarily pseudo-anomaly generator due to its statistical properties and simplicity. Gaussian noise, characterized by its mean and variance, serves as an ideal candidate for generating random deviations that can mimic unexpected patterns in normal data.
@@ C.1. Theoretical Rationale for Gaussian Noise
C.2. Validation Through Total Variation Distance (TVD)
To validate the suitability of Gaussian noise as a pseudo-anomaly generator, we conduct an empirical analysis using the Total Variation Distance (TVD) metric. TVD is a standard measure used in probability theory to quantify the dissimilarity between two distributions. We calculate TVD to evaluate how closely Gaussian noise could mimic true anomalies in both spatial and intensity domains.
@@ C.2. Validation Through Total Variation Distance (TVD)
We conduct experiments on our dataset using both generated Gaussian noise and real anomalies. presents a comparison of the TVD values for different classes. The results indicate that the Gaussian distribution closely matches the distribution of real normal and anomalies across a range of classes. Specifically, it can be seen in  fig. 5.
Table 9. Performance comparison across different methods in multi-class tasks on VisA dataset.
Category DRAEM (Zavrtanik et al., 2021) JNLD (Liu et al., 2023) OmniAL (Zhao, 2023) UniAD (You et al., 2022) UniAD(+COAD)
Single Instance Cashew 67.3 / 68.5 82.5 / 94.1 88.6 / 95.0 92.8/98.6 93.7/98.6
Chewing gum 90.0 / 92.7 96.0 / 98.9 96.4 / 99.0 96.3/98.8 96.6/99.0
Fryum 86.2 / 83.2 91.9 / 90.0 94.6 / 92.1 83.0/95.9 87.9/97.3
Pipe fryum 87.1 / 72.3 87.5 / 92.5 86.1 / 98.2 94.7/98.9 97.8/98.4
Multiple Instances Candle 70.2 / 82.6 85.4 / 94.5 86.8 / 95.8 94.1/98.5 95.8/98.7
Capsules 89.6 / 96.6 91.4 / 99.6 90.6 / 99.4 55.6/88.7 78.3/97.1
Macaroni 1 68.6 / 89.8 90.5 / 93.3 92.6 / 98.6 79.9/97.4 87.2/98.2
Macaroni 2 60.3 / 83.2 71.3 / 92.1 75.2 / 97.9 71.6/95.2 80.4/97.6
Complex Structure PCB1 83.9 / 94.0 82.9 / 98.0 77.7 / 97.6 92.8/93.3 96.9/97.1
PCB2 81.7 / 94.1 79.1 / 95.0 81.0 / 93.9 87.8/93.9 94.1/95.2
PCB3 87.7 / 94.1 90.1 / 98.5 88.1 / 94.7 78.6/97.3 84.5/98.0
PCB4 87.1 / 72.3 96.2 / 97.5 95.3 / 97.1 98.8/94.9 98.7/94.7
Avg. 80.0/85.3 87.1/95.3 87.8/96.6 85.5/95.9 91.0/97.5
C.3. Experimental Results and Qualitative Study
@@
To further validate the results, we present detailed histograms for the anomaly detection scores obtained across 15 individual categories as well as the overall dataset. These histograms are generated under the condition where the Anomaly Rate Quotient (ARQ) is set to 0.006, using RD++ as our baseline framework. Each histogram in  figs. 6 and 7 provides a clear visualization of the distribution of detection scores, overlaid with a Gaussian fitting curve for enhanced interpretability. This fitting curve highlights the distribution, allowing for a more effective comparison of the model’s performance across different categories and the overall dataset, further solidifying the reliability of our evaluation metrics.
C.4. Future Directions with GLASS-like Pseudo-Anomaly Generators
While Gaussian noise has proven effective as a preliminary pseudo-anomaly generator, future work could explore more advanced methods inspired by the GLASS approach (Chen et al., 2024). The GLASS method introduces a structured way to generate pseudo-anomalies by leveraging geometric transformations and localized pattern synthesis, providing greater control over anomaly characteristics.
Appendix D Additional Experimental Results on VisA Dataset
In this section, we provide additional quantitative results for the multi-class anomaly detection task evaluated on the VisA dataset, which complements the results presented in the main paper. Detailed performance comparisons across different methods are summarized in Table 9.
Appendix E Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
To further validate the effectiveness of our methods, we provide a qualitative comparison of anomaly detection results across 15 distinct categories using baselines such as RD, RD++, UniAD, DiAD, and the proposed COAD framework in  figs. 8, 9, 10 and 11. Each subfigure demonstrates: (from left to right) the original image, the ground truth anomaly mask, the result from the baseline method, and the result after applying COAD.
The comparison shows that COAD effectively reduces both false positive rates (FPR) and false negative rates (FNR), accurately capturing anomaly regions while minimizing misclassifications. These results are consistent with the improvements reflected in the earlier quantitative data in  tables 4 and 5, where COAD consistently outperforms its counterparts in image-level and pixel-level AUROC metrics, highlighting COAD’s capability to outperform the baselines.
@@ Appendix E Qualitative Comparison of Anomaly Detection between Different Baselines and COAD
Figure 6. Histograms for anomaly scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
@@
Figure 7. Histograms for normal scores of 15 individual categories. Each histogram includes Gaussian fitting curves for better visualization.
@@
Figure 8. Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using the Baseline Method RD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 9. Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method RD++ and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 10. Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method UniAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
@@
Figure 11. Qualitative Comparison of Anomaly Detection Results Across 15 Categories Using a Baseline Method DiAD and COAD. Each subfigure shows: (from left to right) the original image, ground truth, result from the baseline method, and the result after applying COAD.
Appendix F Future Work
In future work, we plan to refine our pseudo-anomaly generators by incorporating advanced methods in  section C.4, enabling the generation of more structured and domain-relevant pseudo-anomalies.